# One size doesn’t fit all: Predicting the Number of Examples for In-Context Learning

### v2

@@ 4.1 Research Questions and Datasets
  • RQ-3: Does AICL generalise well across different LLMs or datasets?
@@
As per the setup of Ma et al. (2023); Zhang et al. (2022), we conducted our experiments on four standard text classification datasets, namely SST2, TREC Li and Roth (2002), CoLA and RTE. Except TREC, the other three datasets are part of the GLUE benchmark Wang et al. (2018) exhibiting diversity in: i) domain, e.g., from movie reviews to news, ii) number of classes, ranging from 2 to 6, iii) average length of the documents ranging from sentences (e.g., 6.7 words on an average in CoLA) to relatively large text (e.g., 52.4 words on an average in RTE). This diversity in the characteristics of the respective datasets (see Table 1) thus allows provision to assess the generalizability of our proposed approach. More details on each dataset is as follows.
@@
  • TREC: The task is that of open-domain question classification of fact-based questions categorized into a total of 6 types Ma et al. (2023), constituting $5500$ labeled questions in the training set and $500$ in the test set.
@@
  • RTE: The Recognizing Textual Entailment (RTE) dataset comes from a series of annual textual entailment challenges. The authors of the GLUE benchmark Wang et al. (2018) combined the data from RTE1, RTE2, RTE3 and RTE5.
Llama-2-7b Phi-2 Llama-2-13b Method $k$ Prec Rec F1 $k$ Prec Rec F1 $k$ Prec Rec F1 SST2 0-shot 0 .7227 .6123 .5589 0 .8988 .8919 .8914 0 .7341 .6443 .5891 FICL 9 .8744 .8687 .8682 10 .9279 .9253 .9252 9 .8912 .8733 .8809 AICL(E) 8.63 .9089 .8946 .8925 6.39 .9307 .9300 .9300 5.73 .9175 .9065 .9028 AICL(E+N) 8.23 .9099 .8964 .8954 5.64 .9350 .9345 .9345 4.89 .9189 .9071 .9034 AICL* 1.27 .9765 .9753 .9753 .17 .9865 .9863 .9863 .84 .9866 .9859 .9848 TREC 0-shot 0 .1697 .1687 .0100 0 .6289 .4273 .3526 0 .3212 .3685 .3016 FICL 9 .7388 .7324 .6608 8 .6842 .6582 .6192 9 .7612 .7922 .7071 AICL(E) 9.60 .7654 .8049 .7325 4.72 .7146 .7549 .7196 5.51 .7659 .8218 .7383 AICL(E+N) 9.58 .7682 .8075 .7357 5.57 .7254 .7673 .7291 4.93 .7737 .8288 .7532 AICL* 3.35 .8496 .9334 .8616 1.27 .9337 .9294 .9313 2.25 .9513 .9367 .9413 CoLA 0-shot 0 .5585 .5323 .3699 0 .4315 .4899 .2558 0 .6474 .4321 .3455 FICL 10 .6240 .6415 .6008 9 .7071 .6306 .6433 8 .7167 .6338 .6472 AICL(E) 8.32 .6556 .6741 .6582 7.73 .7289 .6471 .6601 2.33 .7412 .6447 .6667 AICL(E+N) 7.43 .6580 .6765 .6616 7.93 .7392 .6486 .6613 1.99 .7432 .6558 .6714 AICL* 2.42 .9417 .9281 .9299 .69 .9663 .9232 .9413 .78 .9408 .9466 .9483 RTE 0-shot 0 .4913 .4992 .3985 0 .6741 .6741 .6741 0 .5345 .5444 .4403 FICL 10 .6095 .6051 .5967 1 .7347 .7239 .7240 4 .6287 .6233 .6214 AICL(E) 8.57 .6304 .6277 .6227 2.44 .7536 .7412 .7415 5.16 .7592 .6441 .6700 AICL(E+N) 8.32 .6342 .6311 .6252 1.15 .7551 .7465 .7471 3.68 .7631 .6485 .6738 AICL* .95 .9783 .9783 .9783 .81 .9296 .9214 .9234 .65 .9734 .9733 .9733
@@ 4.2 Methods Investigated
We compare our proposed AICL methodology with the following baselines:
@@
  • FICL Liu et al. (2022a) (Fixed ICL): This refers to the standard method of applying a static (fixed) number of localized (semantically similar) examples as input. The number of examples to use in FICL is determined by optimizing $k\in\{1,\ldots,10\}$ by means of grid search on a validation set, which comprises $10\%$ of the training set (the same training set is also used in AICL).
@@
  • AICL with Embedding-only - AICL(E): This is an ablation study, where the number of examples for each data instance is only a function of its textual content - more specifically its SBERT Reimers and Gurevych (2019) embedding (see Equation 4).
@@
As part of the analysis, we also report the performance of AICL under an ideal setting, i.e., when one can choose the best prediction from among a set of predictions obtained with different values of $k$ in $k$-shot LLM-based inference by making use of the test set labels. Despite being non-pragmatic, this method provides an upper bound on the effectiveness that could, in principle, be obtained with the AICL framework. To avoid unfair comparisons of AICL* with AICL, we gray out the AICL* results indicating that these results are only to be used as an apex-line.
@@
(a) SST2
(b) TREC
(c) CoLA
(d) RTE
Figure 3: Macro-averaged F1 scores of FICL with different number of examples on the test splits of the different datasets. We also include the AICL results as a point for comparison. Since AICL does not depend on a fixed value of $k$, these results appear as a horizontal line. These plots demonstrate that AICL can be applied on any dataset without requiring to optimize any hyper-parameter (e.g., $k$ in FICL).
@@ 4.3 Model and hyper-parameter settings
To analyze the variations that may be caused due to this choice of LLMs, we conduct our experiments on three different LLMs - two each from the same family with different sizes, and one from a different family. The objective is to analyze the effects of variations in AICL’s performance corresponding to the size of a model from the same family, or variations in characteristics in models across different families. In particular, as instances of LLMs from the same family we use Meta’s Llama-2-7b and Llama-2-13b Touvron et al. (2023), whereas as another LLM from a different family, we use Microsoft’s Phi-2 Li et al. (2023b). Details on these models are as follows: Llama-2 is a collection of open-source pretrained and fine-tuned text models with a maximum context length of 4096 tokens (we use the 7B and the 13B parameter models). Phi-2 Li et al. (2023b) is a transformer with 2.7 billion parameters. It is trained on documents comprising Python code, StackOverflow, websites and natural language text generated by GPT-3.5. Phi-2 with less than 3B parameters has been shown to yield comparable performance with those achieved with much larger number of parameter, e.g., the 13B variant of Llama-2. The maximum input length of Phi-2 is 2048 tokens.
@@ 4.4 Results
Among the two variants of AICL, it can be seen from Table 2 that the use of neighbor class labels (the ‘E+N’ variant) yields better results than the ablated (‘E’ variant), which means that RQ-2 is answered in affirmative. In relation to RQ-3, Table 2 shows that the improvements in AICL with respect to FICL is consistent across different datasets and also across different LLMs.
@@
A solution to reduce training time is to use a small subsample of the training data to learn the parameters of Equation 6. Figure 4 shows the variations in F1 scores for different proportions of the training set used to learn $\kappa(\mathbf{x})$. An interesting observation is that AICL is able to outperform the baseline with as small a proportion as 30% (or even less in some cases) of the training set being used to train $\kappa(\mathbf{x})$, which means that RQ-4 is answered in the affirmative.
5 Conclusion and Future work

### v3

@@ 4.1 Research Questions and Datasets
  • RQ-3: Does AICL generalise well across different LLMs and datasets?
@@
As per the setup of [25, 46], we conducted our experiments on four standard text classification datasets, namely SST2, TREC [17], CoLA and RTE. Except TREC, the other three datasets are part of the GLUE benchmark [43] exhibiting diversity in: i) domain, e.g., from movie reviews to news, ii) number of classes, ranging from 2 to 6, iii) average length of the documents ranging from sentences (e.g., 6.7 words on an average in CoLA) to relatively large text (e.g., 52.4 words on an average in RTE). This diversity in the characteristics of the respective datasets (see Table 1) thus allows provision to assess the generalizability of our proposed approach. More details on each dataset is as follows.
@@
  • TREC: The task is that of open-domain question classification of fact-based questions categorized into a total of 6 types [25], constituting $5500$ labeled questions in the training set and $500$ in the test set.
@@
  • RTE: The Recognizing Textual Entailment (RTE) dataset comes from a series of annual textual entailment challenges. The authors of the GLUE benchmark [43] combined the data from RTE1, RTE2, RTE3 and RTE5.
@@ 4.2 Methods Investigated
We compare our proposed approach [^2] Code available at https://github.com/ManishChandra12/adaptiveICL with the following baselines:
@@
  • FICL [19] (Fixed ICL): This refers to the standard method of applying a static (fixed) number of localized (semantically similar) examples as input. The number of examples to use in FICL is determined by optimizing $k\in\{1,\ldots,10\}$ by means of grid search on a validation set, which comprises $10\%$ of the training set (the same training set is also used in AICL).
@@
  • AICL with Embedding-only - AICL(E): This is an ablation study, where the number of examples for each data instance is only a function of its textual content - more specifically its SBERT [36] embedding (see Equation 4).
@@
We also analyze the performance of AICL under an ideal setting, i.e., when one can choose the best prediction from among a set of predictions obtained with different values of $k$ in $k$-shot LLM-based inference by making use of the test set labels. Despite being non-pragmatic, this method provides an upper bound on the effectiveness that could, in principle, be obtained with the AICL framework. To avoid unfair comparisons of AICL* with AICL, we gray out the AICL* results indicating that these results are only to be used as an apex-line.
@@ 4.3 Model and hyper-parameter settings
To analyze the variations that may be caused due to this choice of LLMs, we conduct our experiments on three different LLMs - two each from the same family with different sizes, and one from a different family. The objective is to analyze the effects of variations in AICL’s performance corresponding to the size of a model from the same family, or variations in characteristics in models across different families. In particular, as instances of LLMs from the same family we use Meta’s Llama-2-7b and Llama-2-13b [42], whereas as another LLM from a different family, we use Microsoft’s Phi-2 [18]. Details on these models are as follows: Llama-2 is a collection of open-source pretrained and fine-tuned text models with a maximum context length of 4096 tokens (we use the 7B and the 13B parameter models). Phi-2 [18] is a transformer with 2.7 billion parameters. It is trained on documents comprising Python code, StackOverflow, websites and natural language text generated by GPT-3.5. Phi-2 with less than 3B parameters has been shown to yield comparable performance with those achieved with much larger number of parameter, e.g., the 13B variant of Llama-2. The maximum input length of Phi-2 is 2048 tokens.
@@
Llama-2-7b Phi-2 Llama-2-13b Method $k$ Prec Rec F1 $k$ Prec Rec F1 $k$ Prec Rec F1 SST2 0-shot 00 .7227 .6123 .5589 00 .8988 .8919 .8914 00 .7341 .6443 .5891 FICL $9$8744 .8687 .8682 $10$9279 .9253 .9252 $9$8912 .8733 .8809 AICL(E) $8.63$9089 .8946 .8925 $6.39$9307 .9300 .9300 $5.73$9175 .9065 .9028 AICL(E+N) $8.23$9099 .8964 .8954 $5.64$9350 .9345 .9345 $4.89$9189 .9071 .9034 AICL* $1.27$9765 .9753 .9753 $0.17$9865 .9863 .9863 $0.84$9866 .9859 .9848 TREC 0-shot 00 .1697 .1687 .0100 00 .6289 .4273 .3526 00 .3212 .3685 .3016 FICL $9$7388 .7324 .6608 $8$6842 .6582 .6192 $9$7612 .7922 .7071 AICL(E) $9.60$7654 .8049 .7325 $4.72$7146 .7549 .7196 $5.51$7659 .8218 .7383 AICL(E+N) $9.58$7682 .8075 .7357 $5.57$7254 .7673 .7291 $4.93$7737 .8288 .7532 AICL* $3.35$8496 .9334 .8616 $1.27$9337 .9294 .9313 $2.25$9513 .9367 .9413 CoLA 0-shot 00 .5585 .5323 .3699 00 .4315 .4899 .2558 00 .6474 .4321 .3455 FICL $10$6240 .6415 .6008 $9$7071 .6306 .6433 $8$7167 .6338 .6472 AICL(E) $8.32$6556 .6741 .6582 $7.73$7289 .6471 .6601 $2.33$7412 .6447 .6667 AICL(E+N) $7.43$6580 .6765 .6616 $7.93$7392 .6486 .6613 $1.99$7432 .6558 .6714 AICL* $2.42$9417 .9281 .9299 $0.69$9663 .9232 .9413 $0.78$9408 .9466 .9483 RTE 0-shot 00 .4913 .4992 .3985 00 .6741 .6741 .6741 00 .5345 .5444 .4403 FICL $10$6095 .6051 .5967 $1$7347 .7239 .7240 $4$6287 .6233 .6214 AICL(E) $8.57$6304 .6277 .6227 $2.44$7536 .7412 .7415 $5.16$7592 .6441 .6700 AICL(E+N) $8.32$6342 .6311 .6252 $1.15$7551 .7465 .7471 $3.68$7631 .6485 .6738 AICL* $0.95$9783 .9783 .9783 $0.81$9296 .9214 .9234 $0.65$9734 .9733 .9733
@@ 4.4 Results
(a) SST2
(b) TREC
(c) CoLA
(d) RTE
Figure 3: FICL (macro-averaged) F-scores for different context sizes on the test splits of the different datasets (AICL results also included for comparison). These plots demonstrate that AICL can be applied on any dataset without requiring to optimize any hyper-parameter (e.g., $k$ in FICL).
@@
Among the two variants of AICL, Table 2 shows that the use of neighbor class labels (the ‘E+N’ variant) yields better results than the ablated (‘E’ variant), which means that RQ-2 is answered in affirmative. In relation to RQ-3, Table 2 shows that the improvements in AICL with respect to FICL is consistent across different datasets and also across different LLMs.
(a) SST2
(b) TREC
(c) CoLA
(d) RTE
@@
A solution to reduce training time is to use a small subsample of the training data to learn the parameters of Equation 6. Figure 4 shows the variations in F1 scores for different proportions of the training set used to learn $\kappa(\mathbf{x})$. While instance selection based sampling methods have been proposed in the literature [4], for this work we simply use uniform sampling and report the results averaged over 5 runs of the experiments.
An interesting observation is that AICL is able to outperform the baseline with as small a proportion as 30% (or even less in some cases) of the training set being used to train $\kappa(\mathbf{x})$, which means that RQ-4 is answered in the affirmative.
Sensitivity to Labels.
We now analyze how AICL behaves in the absence of any labelled training data [27] - a setup commonly known as retrieval-augmented generation (RAG) [9]. A way to simulate this configuration in our experiment setup is to keep only the content and exclude the labels from the demonstrations added to the prompt. Equation 3, in its current form, cannot be used to get all the numbers of examples that work, because of the unavailability of labels. We therefore modify Equation 3 to set $\pazocal{K}_{i}(\mathbf{x})=1$, where $0\leq i\leq M$ denotes the number of examples that lead to the maximum mutual information of the Softmax class posteriors, as obtained with $i$-shot LLM output. The rest of the components of $\pazocal{K}_{i}(\mathbf{x})$ are set to 0, which means that instead of a multi-label classification setting, for the unlabeled setup the $k$ predictor is now a multi-class classifier. Also note that due to the unavailability of class labels, the AICL(E+N) version (Equation 6) cannot be executed in this RAG setup.
RAG setup (w/o Labels) ICL setup (w/ Labels) Dataset 0-shot FICL AICL(E) AICL* FICL AICL(E) AICL* SST2 .8914 .7339 .9119 .9610 .9252 .9300 .9863 TREC .3526 .4287 .4752 .4922 .6192 .7196 .9313 CoLA .2558 .2469 .2679 .7937 .6433 .6601 .9413 RTE .6741 .6144 .6688 .8655 .7240 .7415 .9234
Table 3: A comparison between the F1 scores obtained by AICL (Phi-2) on two different settings - a) w/ labels (same as Table 2), and b) w/o labels, as in a RAG pipeline. Best performing RAG results are bold-faced.
The results of these additional experiments, carried out on the Phi-2 model, are shown in Table 3. From Table 3, it can be seen that absence of labels hurt performance - however, not by large margins, which is consistent with the findings of [27]. AICL still continues to mostly be the best performing method even without the presence of any class labels. The only dataset where AICL does not yield the best performance is RTE, where the best performance is achieved with 0-shot indicating that labels are crucial for this dataset, and hence all RAG approaches fail on this particular dataset. It is encouraging to see, however, that AICL* (the oracle setting of AICL) without labels still continues to be substantially better than FICL (ICL with fixed context sizes).
Concluding Remarks.
@@
{credits}
[…]
The authors have no competing interests to declare that are relevant to the content of this article.